A body travels from A to B at 25m/s and from B to A at 50m/s,then it's average speed will be

Grade:12th pass

1 Answers

Heer Shah
2 years ago
Average speed=total distance travel/Total time taken
Let distance from A to B be x and let it be a straight path then,
Total distance travelled =2x
Time taken from A to B=x/25 
Time taken from B to A=x/50
Total Time taken =x×(1/25+1/50)=3x/50
Average speed =2x/(3x/50)=100/3 m/s